# WildFireCan-MMD: A Multimodal Dataset for Classification of User-Generated Content During Wildfires in Canada

### v3

@@ Appendix F VLMs as Zero-Shot Classifiers:
LLaVA-13B: Also considered an intermediately-sized option, the 13B 4-bit quantized version of LLaVA was used due to hardware constraints. Using a combination of a publicly available Colab notebook  [^7] https://github.com/camenduru/LLaVA-colab, the author of which performed the quantization, and the original code from the open-source LLaVA repository  [^8] https://github.com/haotian-liu/LLaVA, the model was able to be used with Google Colab.
GPT-4o-mini: The largest of the VLMs tested, a paid API service had to be used for querying GPT-4o-mini 999version:GPT-4o-mini-2024-07-18, and it is also the only closed-source model tested. A specified content field in each query contains the system prompt, text and image from the tweet, and user prompt.
For each method, the responses are saved as a file where each line contains the model’s answer supplied as a letter from the options shown in Table D.1. This then allows for the calculation of metrics.
Appendix G Related Architectures
Table G.1 briefly describes the architectures proposed by various authors on the CrisisMMD dataset. The purpose of this table is to demonstrate the literature review used to select the classifier proposed in section 4 and give an idea of the expected accuracy of such a classifier on this task of wildfire social-media data.
@@ Appendix G Related Architectures
Agar. 2020 Agarwal et al. (2020) InceptionV3+RCNN+AttentionFusion 99.00 99.00 97.00
Abav. 2020 Abavisani et al. (2020) DenseNet+BERT+CrossAttentionFusion 89.35 91.82 70.41
Ofli. 2020 Ofli et al. (2020) VGG16+CNN+Concat.Fusion 84.20 78.30 -
Xukun 2020 Li and Caragea (2020) AlexNet+GRU+Concat.Fusion 78.13 - -
Madi. 2021 Madichetty et al. (2021) DenseNet+BERT+AverageFusion 74.85 - -
Kumar 2022 Kumar et al. (2022) VGG16+LSTM+Concat.Fusion 82.00 - -
Pran. 2022 Pranesh (2022) ResNet+RoBERTa+TransformerAttentionFusion 85.10 89.50 -
Sirbu 2022 Sirbu et al. (2022) semi-supervised 82.00 87.20 88.70
Diva. 2022 Divakaran et al. (2022) DenseNet+RoBERTa+Concat.Fusion 89.26 85.87 -
Kotha 2022 Kotha et al. (2022) RegNetY320+BERT+Concat.Fusion 89.63 86.59 -
Khat. 2022 Khattar and Quadri (2022) VGG16+BiLSTM+CrossAttentionFusion 84.08 - -
Madi. 2023 Madichetty et al. (2023) VGG16+RoBERTa+MultiplicativeFusion 90.21 - -
Basit 2023 Basit et al. (2023) DeIT+ALBERT+Concat.Fusion 79.00 80.00 86.00
Koshy 2023 Koshy and Elango (2023) ViT+RoBERTa+BiLSTM+BilinearFusion 97.14 - -
Gupta 2024 Gupta et al. (2024) DenseNet+ELECTRA+CrossAttentionFusion 91.20 93.40 72.20
Shet. 2025 Shetty et al. (2025) ViT+RoBERTa+CrossAttentionFusion 91.25 89.63 68.72
Sirbu 2025 Sîrbu et al. (2025) semi-supervised 91.22 89.55 -
Teng 2025 Teng and Öhman (2025) VGG16+BERT+CrossAttentionFusion 81.00 - -
Dar 2025 Dar et al. (2025) (ResNet50+BERT)+(Graph)+IDEAFusion 98.23 90.13 -
@@
Table G.1: Literature review of deep-learning approaches used by other works on the CrisisMMD dataset.
Appendix H Ablation & Other Experiments
Tables H.1 and H.2 show the different text/image encoders experimented with, using either only text or only images respectively. All experiments here were conducted using the same train/test split, which was earlier identified as being a good split through cross validation experiments on different random seeds. Experiments were done on different final feature representations by using either mean pooling of the last hidden state or, using the CLS token. Also, freezing the first half of the layers as this has been shown to increase accuracy Ingle et al. (2022). All encoders had a linear layer appended for classification.
For text, "roberta-base" performed best when using just the CLS token for classification and freezing the first half of the encoder layers, as shown in Table H.1. Later experimentation found that the unfrozen version performed better in the final classifier. For images, "vit-base-patch16-384" performed best when using just the CLS token for classification without any freezing.
Many of the fusion experiments tested, as well as some architectures from other studies can be seen in Table H.3. The same train/test split was used as before for unimodal classification. The best result, belonging to the selected classifier, is shown in bold.
@@ Appendix H Ablation & Other Experiments
Table H.4 shows the ablation study performed on the selected model. (1) is the result of the selected model, (2) is the result of the model with a simpler classification head consisting of a single linear layer, (3) is the model with the proposed classification head but without the fusion module,
Three-fold cross-validation was performed later to support the result as described in section 4.
Text Model Notes F1-Weighted
’bert-base-cased’ mean-pooling 0.7779
’bert-base-cased’ CLS 0.7793
"bertweet-base" mean-pooling 0.7900
"bertweet-base" CLS 0.7927
"ModernBERT-base" mean-pooling 0.7934
"ModernBERT-base" CLS 0.7862
"roberta-base" robertaforseq 0.7718
"roberta-base" mean-pooling 0.7934
"roberta-base" mean-pooling, half-frozen 0.7940
"roberta-base" mean-pooling, custom vocab 0.7209
"roberta-base" CLS 0.7798
"roberta-base" CLS, half-frozen 0.8008
Table H.1: Text-only results.
Image Model Notes F1-Weighted
DenseNet121 ImageNet weights 0.5733
InceptionV3 ImageNet weights 0.5804
ResNet50 ImageNet weights 0.5841
VGG16 ImageNet weights 0.5993
"facebook/deit-base-patch16-384" CLS 0.6260
"google/vit-base-patch16-384" half-frozen, CLS 0.6290
"google/vit-base-patch16-384" mean-pooling 0.6404
"google/vit-base-patch16-384" CLS 0.6419
Table H.2: Image-only results.
Text Model Image Model Fusion Type Notes F1
RoBERTa+BiLSTM ViT BiLSTM+BiLinear Koshy’s model Koshy and Elango (2023) 0.6149
CLIP CLIP "openai/clip-vit-base-patch32" 0.7072
VILT VILT "dandelin/vilt-b32-finetuned-nlvr2" 0.7638
RoBERTa ViT Transformer RoBERTa/ViT fully frozen 0.6164
RoBERTa x2 ViT Concatenative extract text from images with OCR 0.8102
RoBERTa ViT Concatenative roberta:mean, vit:cls 0.8028
Table H.3: Multimodal results.
@@
Table H.4: Ablation of model components.
Appendix I Structure of Custom Classifier
Figure I.1 shows the architecture of the selected approach. This model is described in detail in Section 4.
Figure I.1: Selected dual-encoder architecture utilizing transformer fusion module.
Appendix J Confusion Matrices
For deeper insights into the classifiers performances, the confusion matrices are shown in Figures J.1 and J.2.
@@ Appendix I Structure of Custom Classifier
Figure J.1: Confusion matricies for selected classifier over the different train/test splits.
@@
Figure J.2: Confusion matrices of VLMs.
Appendix K Queries Used for Additional Data Collection
@@
Appendix L Unlabeled Dataset
Figure 2 shows the distribution of post origins. A word cloud of the text data of the unlabeled dataset collected in Section 7, and also the frequency distribution of the posts by word count, are shown in Figure L.1.
@@
Figure L.1: (a) Word cloud of the text from the unlabeled tweets and (b) the frequency distribution of their word counts.

### v4

@@ Appendix E VLMs as Zero-Shot Classifiers:
LLaVA-v1.5-13B: Also considered an intermediately-sized option, the 13B 4-bit quantized version of LLaVA was used due to hardware constraints. Using a combination of a publicly available Colab notebook  [^7] https://github.com/camenduru/LLaVA-colab, the author of which performed the quantization, and the original code from the open-source LLaVA repository  [^8] https://github.com/haotian-liu/LLaVA, the model was able to be used with Google Colab.
GPT-4o-mini: The largest of the VLMs tested, a paid API service had to be used for querying GPT-4o-mini 999version:GPT-4o-mini-2024-07-18, and it is also the only closed-source model tested.
For each method, the responses are saved as a file where each line contains the model’s answer supplied as a letter from the options shown in Table C.1. This then allows for the calculation of metrics.
Appendix F Related Architectures
Table F.1 briefly describes the architectures proposed by various authors for the CrisisMMD dataset. The purpose of this table is to demonstrate the literature review used to select the classifier proposed in section 4 and give an idea of the expected accuracy of such a classifier on this task of wildfire social-media data.
@@ Appendix F Related Architectures
Agar. 2020 Agarwal2020 InceptionV3+RCNN+AttentionFusion 99.00 99.00 97.00
Abav. 2020 Abavisani2020 DenseNet+BERT+CrossAttentionFusion 89.35 91.82 70.41
Ofli. 2020 Ofli2020 VGG16+CNN+Concat.Fusion 84.20 78.30 -
Xukun 2020 xukun2020 AlexNet+GRU+Concat.Fusion 78.13 - -
Madi. 2021 Madichetty2021 DenseNet+BERT+AverageFusion 74.85 - -
Kumar 2022 kumar2022 VGG16+LSTM+Concat.Fusion 82.00 - -
Pran. 2022 pranesh2022 ResNet+RoBERTa+TransformerAttentionFusion 85.10 89.50 -
Sirbu 2022 sirbu2022 semi-supervised 82.00 87.20 88.70
Diva. 2022 Divakaran2022 DenseNet+RoBERTa+Concat.Fusion 89.26 85.87 -
Kotha 2022 kotha2022 RegNetY320+BERT+Concat.Fusion 89.63 86.59 -
Khat. 2022 Khattar2022 VGG16+BiLSTM+CrossAttentionFusion 84.08 - -
Madi. 2023 Madichetty2023 VGG16+RoBERTa+MultiplicativeFusion 90.21 - -
Basit 2023 Basit2023 DeIT+ALBERT+Concat.Fusion 79.00 80.00 86.00
Koshy 2023 Koshy2023 ViT+RoBERTa+BiLSTM+BilinearFusion 97.14 - -
Gupta 2024 Gupta2024 DenseNet+ELECTRA+CrossAttentionFusion 91.20 93.40 72.20
Shet. 2025 shetty2025 ViT+RoBERTa+CrossAttentionFusion 91.25 89.63 68.72
Sirbu 2025 sirbu2025 semi-supervised 91.22 89.55 -
Teng 2025 teng2025 VGG16+BERT+CrossAttentionFusion 81.00 - -
Dar 2025 Dar2025 (ResNet50+BERT)+(Graph)+IDEAFusion 98.23 90.13 -
@@
Table F.1: Literature review of deep-learning approaches used by other works on the CrisisMMD dataset.
Appendix G Ablation & Other Experiments
@@
Table G.1 shows the ablation study performed on the selected model. (1) is the result of the selected model, (2) is the result of the model with a simpler classification head consisting of a single linear layer, (3) is the model with the proposed classification head but without the fusion module. Two other random splits were later used to support the result described in section 4.
@@
Table G.1: Ablation of model components.
Appendix H Structure of Custom Classifier
Figure H.1 shows the architecture of the selected approach. This model is described in detail in Section 4.
Figure H.1: Selected dual-encoder architecture utilizing transformer fusion module.
Appendix I Confusion Matrices
For deeper insights into the classifiers performances, the confusion matrices are shown in Figures I.1 and I.2.
@@ Appendix I Confusion Matrices
Figure I.1: Confusion matricies for selected classifier over the different train/test splits.
@@
Figure I.2: Confusion matrices of VLMs.
Appendix J Queries Used for Additional Data Collection
@@
Appendix K Unlabeled Dataset
Figure 2 shows the distribution of post origins. A word cloud of the text data of the unlabeled dataset collected in Section 7, and also the frequency distribution of the posts by word count, are shown in Figure K.1.
@@
Figure K.1: (a) Word cloud of the text from the unlabeled tweets and (b) the frequency distribution of their word counts.